OpenStudy (anonymous):

solve the proportion x+6/x=5/4

OpenStudy (anonymous):

x+6/x=5/4 or (x+6)/x = 5/4 which is it ?

OpenStudy (anonymous):

x+6/x=5/4

OpenStudy (anonymous):

so only the 6 is divided by x right?

OpenStudy (anonymous):

Sorry it is (x+6)/x=5/4